# Cavity modification of magnetoplasmon mode through coupling with intersubband polaritons

### v1

@@ IV.3 Generality of Non-Locality
For sufficiently doped quantum wells, the nonlocality behavior only depends on the inhomogeneity of the cavity mode, and can thus be recovered in other types of heterostructure, and for different cavity geometries. To demonstrate this, in Figure 4 we show the case of ISB polaritons in a cavity containing parabolic quantum wells [11]. Here, the ISBT is still coupled to the TM mode of the cavity, but the continuously-graded material composition of the quantum well allows for a lower frequency ISB transition, and therefore larger ISB polariton coupling strength. To match the TM mode frequency to the ISB frequency, the cavity dimensions are slightly different from the square well case, with no GaAs buffer region (Fig. 4a). Despite the altered cavity dimensions, we still observe significant inhomogeneity in the fields, particularly at the interface between the metallic cavity plate and the gap. In addition, the lack of a buffer region means that the electrons in the quantum wells also experience the fields at the grating edges, which are highly inhomogeneous.
@@
As for the previous samples, Figure 4b shows the measured spectra when a magnetic field is applied, tuning the MP resonance across the ISB polaritons. Here, interaction of the MP with both polariton branches is visible thanks to the closer frequency alignment between the TM cavity mode and ISB, although the spectral features are slightly less clear due to the broadened polariton branches. For both polariton branches, these features include a cavity-induced broadening around the crossing point, a blue-shift of the MP from its bare frequency $\omega_{B}$ indicated as a black line, and an eventual red-shift of the polariton branches at strong $B$ fields, as seen in the square well case in Figure 3.
@@
By demonstrating the effect of Coulomb interaction in different cavities with two different heterostructures, we highlight that the level of nonlocality can be tailored through cavity design, independently from the quantum well system. To further support this, in Appendix D we show the calculated spectra for both the cavities in Fig. 3 and Fig. 4 without including Coulomb interaction, demonstrating how the experimental spectra is clearly modified in the experiment by nonlocality. Moreover, by further increasing the coupling strength either through different cavity design or increasing the electron density, the Coulomb interaction and thereby nonlocality can be further enhanced.
@@ V Conclusion
We find that the MP–cavity coupling in spatially inhomogeneous field configurations results in significant new features, including spectral shifts attributable to Coulomb interaction. This breakdown of translational invariance - induced purely by the spatial profile of the cavity field - demonstrates a method for altering materials through cavity coupling that arises not from vacuum fluctuations but simply from the cavity field spatial profile. Furthermore, the cavity-induced Coulomb interactions can be activated on demand via magnetic field tuning and are controlled through the cavity design.
These findings broaden the understanding of light–matter interactions in the strong coupling regime and provide a versatile route for tailoring electronic excitations in solid-state systems. Beyond this, the ability to manipulate polariton properties dynamically through cavity mode selection and magnetic field tuning opens new possibilities for tunable polaritonic devices, where the tripartite interplay can be also exploited to study the cavity protection phenomena or the quantum vacuum of ultrastrongly coupled systems [46, 47].
@@ A.3 Cyclotron polarization and non-locality
$\partial_{t}^{2}\textbf{P}(\textbf{r})+\omega_{B}^{2}\textbf{P}(\textbf{r})=\omega_{P}L_{c}\epsilon\textbf{E}(\textbf{r},0).$ (38)
@@ Appendix D Comparison to simulation without the longitudinal Coulomb contribution
In this last section, we present a simulation example with the same parameters as described above and used in the main text, but where we artificially remove the shift induced by Coulomb effects, setting $\bar{\omega}_{B}(k)=\omega_{B}$. The result is in Fig. 6. Without the inclusion of the k-dependent shift due to the longitudinal Coulomb interaction, we can see much better the Rabi splitting between the MP and the polariton lines, which is perfectly symmetric. This type of plot does not correctly match the experimental result, as one can directly check by comparison with Fig. 3 and 4.
Figure 6: (a-c) Same simulation as in Fig. 3. (b-d) Same simulation as in Fig. 4. In both cases $\zeta_{k}=0$.

### v2

@@ IV.4 Generality of Non-Locality
For sufficiently doped quantum wells, the non-locality behavior only depends on the inhomogeneity of the cavity mode, and can thus be recovered in other types of heterostructure, and for different cavity geometries. To demonstrate this, in Figure 4 we show the case of ISB polaritons in a cavity containing parabolic quantum wells [11]. Here, the ISBT is still coupled to the TM mode of the cavity, but the continuously-graded material composition of the quantum well allows for a lower frequency ISB transition, and therefore larger ISB polariton coupling strength. To match the TM mode frequency to the ISB frequency, the cavity dimensions are slightly different from the rectangular well case, with no GaAs buffer region (Fig. 4a). Despite the altered cavity dimensions, we still observe significant inhomogeneity in the fields, particularly at the interface between the metallic cavity plate and the gap. In addition, the lack of a buffer region means that the electrons in the quantum wells also experience the fields at the grating edges, which are highly inhomogeneous.
@@
As for the previous samples, Figure 4b shows the measured spectra when a magnetic field is applied, tuning the MP resonance across the ISB polaritons. Here, interaction of the MP with both polariton branches is visible thanks to the closer frequency alignment between the TM cavity mode and ISB, although the spectral features are slightly less clear due to the broadened polariton branches. For both polariton branches, these features include a cavity-induced broadening around the crossing point, a blue-shift of the MP from its bare frequency $\omega_{B}$ indicated as a black line as seen in the rectangular well case in Figure 3.
@@
By demonstrating the effect of Coulomb interaction in different cavities with two different heterostructures, we highlight that the level of non-locality can be tailored through cavity design, independently from the quantum well system. To further support this, in Appendix D we show the calculated spectra for both the cavities in Fig. 3 and Fig. 4 without including Coulomb interaction, demonstrating how the experimental spectra is clearly modified in the experiment by non-locality. Moreover, by further increasing the coupling strength either through different cavity design or increasing the electron density, the Coulomb interaction and thereby non-locality can be further enhanced.
@@ V Conclusion
We find that the MP–cavity coupling in a spatially inhomogeneous field results in significant new features, including spectral shifts attributable to Coulomb interaction. Furthermore, the breakdown of translational invariance - induced purely by the spatial profile of the cavity field - demonstrates a method for altering materials through cavity coupling that arises not from vacuum fluctuations but simply from the cavity field spatial profile. The cavity-induced Coulomb interactions can be activated on demand via magnetic field tuning and are controlled through the cavity design.
These findings demonstrate that, in order to accurately model light-matter coupled systems with electronic transitions and inhomogeneous electric fields, it is crucial to include the influence of the field inhomogeneity on the Coulomb interactions in the 2DEG. Beyond this, the ability to manipulate polariton properties dynamically through cavity mode selection and magnetic field tuning opens new possibilities for tunable polaritonic devices, where the tripartite interplay can be also exploited to study the cavity protection phenomena or the quantum vacuum of ultrastrongly coupled systems [49, 48].
@@ Acknowledgements.
Data Availability
The data supporting the findings of this article are openly available at [50].
@@ A.3 Cyclotron polarization and non-locality
$\partial_{t}^{2}\textbf{P}(\textbf{r})+\omega_{B}^{2}\textbf{P}(\textbf{r})=\omega_{P}^{2}L_{c}\epsilon\textbf{E}(\textbf{r},0).$ (44)
@@ Appendix B Three-modes cavity theory
To better visualize the effect of the non-locality (as in Fig. 3b), we can define an MP trajectory in the $B$-field, that interpolates between the bare cyclotron resonance $\omega_{B}$ and the brightest k-mode of the MP interacting with the TM cavity mode. Indeed, looking at Eq. (66), we realize that $k_{1}$ is the brightest mode when the MP is resonant with the TM cavity mode, while out of resonance the dominant MP mode is at $k=0$. We then introduce a $B$-dependent $k$ wavevector
$k(B)=k_{1}\frac{\gamma_{\rm TM}^{2}}{\left(\omega_{\rm TM}-\bar{\omega}_{B}(k_{1})\right)^{2}+\gamma_{\rm TM}^{2}}.$ (71)
The red line MP trajectory in Fig. 3b is then simply given by $\omega_{\rm MP}(B)=\bar{\omega}_{B}(k(B))$, calculated on the parameters in the next Section C. Naturally, this $\omega_{\rm MP}(B)$ trajectory is not rigorous, and the interpolation can be done with any other function rather than the Lorentzian used here. However, we find that it correctly visualizes the MP behavior.
@@ Appendix D Comparison to simulation without the longitudinal Coulomb contribution
In this section we further demonstrate the effect of non-locality by comparing our calculated spectra in Figures 3 and 4 to the case where we artificially remove the shift induced by Coulomb effects, setting $\bar{\omega}_{B}(k)=\omega_{B}$. In Fig. 6 we show this comparison for the RQW case (Fig. 3). Without the inclusion of the k-dependent shift due to the longitudinal Coulomb interaction, Rabi splitting between the MP and the polariton lines is more evident and the interaction is perfectly symmetric about the cyclotron resonance frequency. For completeness, in Fig. 7 we also show the calculated spectra without non-locality for the PQW case (for comparison to Fig. 4d.
Figure 6: a) Same simulation as in Fig. 3. Red solid line is a guide for the eye obtained by interpolating between the bare cyclotron frequency and the brightest, blue-shifted, cyclotron k-mode (see Appendix B for further details). The green shaded region represents a collection of the various trajectories $\bar{\omega}_{B}(k)$ for increasing values of $k\in(0,\infty)$; the opacity of each line is given $|w_{\rm TM,k}|^{2}\gamma^{2}/((\omega_{\rm TM}-\bar{\omega}_{B}(k))^{2}+\gamma^{2})$ accordingly to Eq. (66). b) Same simulation as in (a) but with $\bar{\omega}_{B}(k)=\omega_{B}$. c) Individual spectra of TM LP for three values of the magnetic field extracted from (a)-top and (b)-bottom, as in Fig. 3.
Figure 7: a) Same simulation as in Fig. 4, with $\bar{\omega}_{B}(k)=\omega_{B}$. b) Individual spectra of lower (left) and upper (right) polaritons for three values of the magnetic field extracted from (a), as in Fig. 4.
Appendix E Hamiltonian formulation
Here we consider the finalized coupled modes equations in Eqs. (53)-(54)-(55) and we recast this system into the Hamiltonian framework. For simplicity here we omit the tilde $\tilde{\cdot}$ from all the variables.
First we introduce the two auxiliary parameters $\varepsilon_{\rm aux},M_{\rm aux}$, defined such that
$N_{\rm qw}\omega_{P}^{2}=\frac{1}{\varepsilon_{\rm aux}M_{\rm aux}}.$ (77)
Then we introduce a new variable for the electromagnetic field
$\mathcal{E}_{\lambda}=\varepsilon_{\rm aux}\partial_{t}A_{\lambda},$ (78)
and for the polarization
$P_{\lambda}=\sum_{k}\left(\tilde{w}_{\lambda,k}\tilde{P}_{k}^{\rm mp}+\tilde{u}_{\lambda,k}\tilde{P}_{k}^{\rm isb}\right).$ (79)
Then Eq. (53) is rewritten as
$\partial_{t}^{2}A_{\lambda}=-\omega_{\lambda}^{2}A_{\lambda}-\partial_{t}\frac{P_{\lambda}}{\varepsilon_{\lambda}}.$ (80)
Introducing the canonical variable
$D_{\lambda}=\varepsilon_{\rm aux}\partial_{t}A_{\lambda}+P_{\lambda}$ (81)
so we rewrite the equation as the first Hamilton equation
$\partial_{t}D_{\lambda}=-\omega_{\lambda}^{2}A_{\lambda}=-\frac{\partial H_{\mathcal{E}}}{\partial A_{\lambda}},$ (82)
where the electric field Hamiltonian is introduced as
$H_{\mathcal{E}}=\sum_{\lambda}\frac{\left(D_{\lambda}-P_{\lambda}\right)^{2}}{2\varepsilon_{\rm aux}}+\frac{\epsilon_{\rm aux}\omega_{\lambda}^{2}}{2}A_{\lambda}^{2}.$ (83)
By construction also the second Hamilton equation is satisfied
$\partial_{t}A_{\lambda}=\frac{\partial H_{\mathcal{E}}}{\partial D_{\lambda}}=\frac{D_{\lambda}-P_{\lambda}}{\varepsilon_{\rm aux}},$ (84)
due to the definition given in Eq. (81).
In a similar way we derive the Hamiltonian for the MP and ISB polarization equations. Let’s introduce the canonical moment of the MP/ISB degree of freedom
$\Pi_{k}^{\rm mp/isb}=M_{\rm aux}\partial_{t}P_{k}^{\rm mp/isb}.$ (85)
Eqs. (54)-(55) are then rewritten as a first Hamilton equation
$\begin{split}\partial_{t}\Pi_{k}^{\rm mp/isb}&=-\frac{\partial H_{\rm mp/isb}}{\partial P_{k}^{\rm mp/isb}}-\frac{\partial H_{\mathcal{E}}}{\partial P_{k}^{\rm mp/isb}}\\ &=-M_{\rm aux}\omega_{\rm mp/isb}^{2}P_{k}^{\rm mp/isb}+\sum_{\lambda}w_{k\lambda}\frac{D_{\lambda}-P_{\lambda}}{\varepsilon_{\rm aux}},\end{split}$ (86)
where
$H_{\rm mp}=\sum_{k}\frac{\left(\Pi_{k}^{\rm mp}\right)^{2}}{2M_{\rm aux}}+\frac{M_{\rm aux}\bar{\omega}_{B}^{2}(k)}{2}\left(P_{k}^{\rm mp}\right)^{2}$ (87)
$H_{\rm isb}=\sum_{k}\frac{\left(\Pi_{k}^{\rm isb}\right)^{2}}{2M_{\rm aux}}+\frac{M_{\rm aux}\omega_{\rm isb}^{2}}{2}\left(P_{k}^{\rm isb}\right)^{2}.$ (88)
The total Hamiltonian of the system is then
$H=H_{\mathcal{E}}+H_{\rm isb}+H_{\rm mp}.$ (89)
Introducing the complex variables $a_{\lambda}$, $b_{k}$, $c_{k}$, respectively for the cavity, the MP and the ISB, we define
$\begin{split}D_{\lambda}&=-i\sqrt{\frac{\varepsilon_{\rm aux}\omega_{\lambda}}{2}}\left(a_{\lambda}-a_{\lambda}^{{\dagger}}\right),\\ A_{\lambda}&=\sqrt{\frac{1}{2\varepsilon_{\rm aux}\omega_{\lambda}}}\left(a_{\lambda}+a_{\lambda}^{{\dagger}}\right),\end{split}$ (90)
$\begin{split}\Pi_{k}^{\rm mp}&=-i\sqrt{\frac{M_{\rm aux}\omega_{B}}{2}}\left(b_{k}-b_{k}^{{\dagger}}\right),\\ P_{k}^{\rm mp}&=\sqrt{\frac{1}{2M_{\rm aux}\omega_{B}}}\left(b_{k}+b_{k}^{{\dagger}}\right),\end{split}$ (91)
$\begin{split}\Pi_{k}^{\rm isb}&=-i\sqrt{\frac{M_{\rm aux}\omega_{\rm isb}}{2}}\left(c_{k}-c_{k}^{{\dagger}}\right),\\ P_{k}^{\rm isb}&=\sqrt{\frac{1}{2M_{\rm aux}\omega_{\rm isb}}}\left(c_{k}+c_{k}^{{\dagger}}\right).\end{split}$ (92)
The Hamiltonian $H=H_{\mathcal{E}}+H_{\rm mp}+H_{\rm isb}$ becomes
$\begin{split}H_{\mathcal{E}}&=\sum_{\lambda}\omega_{\lambda}a_{\lambda}^{{\dagger}}a_{\lambda}\\ &-i\frac{\sqrt{N_{\rm qw}}\omega_{P}}{2}\sum_{\lambda,k}\left(a_{\lambda}-a_{\lambda}^{{\dagger}}\right)\left[\tilde{w}_{\lambda,k}\left(b_{k}+b_{k}^{{\dagger}}\right)+\tilde{u}_{\lambda,k}\left(c_{k}+c_{k}^{{\dagger}}\right)\right]\\ &+\sum_{\lambda}\frac{N_{\rm qw}\omega_{P}^{2}}{4\omega_{\lambda}}\left[\sum_{k}\tilde{w}_{\lambda,k}\left(b_{k}+b_{k}^{{\dagger}}\right)+\tilde{u}_{\lambda,k}\left(c_{k}+c_{k}^{{\dagger}}\right)\right]^{2},\end{split}$ (93)
$\begin{split}H_{\rm mp}&=\sum_{k}\left[\omega_{B}b^{{\dagger}}_{k}b_{k}+\frac{N_{\rm qw}\omega_{P}^{2}}{4\omega_{B}}\zeta_{k}\left(b_{k}+b_{k}^{{\dagger}}\right)^{2}\right]\\ H_{\rm isb}&=\sum_{k}\omega_{\rm isb}c^{{\dagger}}_{k}c_{k}.\end{split}$ (94)
By reintroducing $\hbar$ and imposing the right commutation relation one has immediately also the quantized three-partite polaritonic system.